### v1

# HumAIne-Chatbot: Real-Time Personalized Conversational AI via Reinforcement Learning

Abstract
Current conversational AI systems often provide generic, one-size-fits-all interactions that overlook individual user characteristics and lack adaptive dialogue management. To address this gap, we introduce HumAIne-chatbot, an AI-driven conversational agent that personalizes responses through a novel user profiling framework. The system is pre-trained on a diverse set of GPT-generated virtual personas to establish a broad prior over user types. During live interactions, an online reinforcement learning agent refines per-user models by combining implicit signals (e.g., typing speed, sentiment, engagement duration) with explicit feedback (e.g., likes and dislikes). This profile dynamically informs the chatbot’s dialogue policy, enabling real-time adaptation of both content and style. To evaluate the system, we conducted controlled experiments with 50 synthetic personas across multiple conversation domains. Results showed consistent improvements in user satisfaction, personalization accuracy, and task achievement when personalization features were enabled. Statistical analysis confirmed significant differences between personalized and non-personalized conditions, with large effect sizes across key metrics. These findings highlight the effectiveness of AI-driven user profiling and provide a strong foundation for future real-world validation.
I Introduction
The Information Age, marked by the proliferation of advanced computational systems and significant data generation, has guided in a transition from Industry 4.0 to Industry 5.0 [1]. Industry 5.0 emphasizes collaboration between humans and robots, incorporating resilience and sustainability goals into industrial practices. While Industry 4.0 focused on technologies such as the Internet of Things (IoT) and big data - with Artificial Intelligence (AI) as its driving force, facilitating the development of innovative tools and methods [2], [3], [4] - Industry 5.0 reintroduces human, environmental, and social considerations into technological advancements. Alongside this shift, the rise of large language models (LLMs) and Generative AI (GenAI) has further enhanced human-machine collaboration.
This evolution has significantly influenced the domain of conversational AI. In recent years, chatbots have become increasingly prevalent across various sectors, transforming how businesses interact with customers and optimizing internal operations. The advent of LLMs like GPT-3 and GPT-4 [5] has enabled chatbots to generate coherent and contextually rich dialogue across diverse topics, expanding the boundaries of conversational AI.
Despite their widespread adoption, current chatbots face significant challenges in delivering a personalized and efficient interaction experience. Two primary issues are particularly notable:
  • Lack of Personalization: Many chatbots today offer generic responses that do not consider individual user characteristics, preferences, or expertise [6].
  • Lack of Personalization metrics: Despite their increasing use, evaluating chatbots remains a challenging research problem, with no universally accepted metric for effectiveness, particularly from user-centric perspectives such as motivation and satisfaction [7], [8]
This paper tackles the challenge of personalized conversational AI by introducing HumAIne-chatbot, an AI-driven chatbot designed to enhance human-machine interactions through adaptive user profiling. The main contributions of this paper are:
  1. 1.
    Developing a user profiling mechanism that integrates implicit and explicit metrics to personalize chatbot responses according to individual user characteristics, preferences, and expertise.
  2. 2.
    Leveraging an innovative virtual personas mechanism to enhance user profiling consistency and robustness.
  3. 3.
    Offering a mechanism for transforming user profiling outputs into actionable prompt refinement, enabling more effective user personalization.
In this context, HumAIne-chatbot leverages AI-driven user profiling to enhance user experience and effectiveness through two personalization methods. First, it enriches prompts to tailor responses according to the user’s knowledge level and expertise, ensuring that the interaction is adaptive to individual user needs.
To achieve these personalization goals, HumAIne-chatbot employs a combination of implicit and explicit metrics to k̈nowẗhe user through AI-based profiling. Implicit metrics include session duration, response time, sentiment analysis, frequency of grammatical mistakes, language complexity, and typing speed. Explicit metrics encompass direct user feedback, profile completeness, and survey-based satisfaction ratings [9, 6]. By utilizing these collected metrics, the AI-based user profiling allows for a personalized interaction experience. Similar approaches have been used in previous studies to gather detailed measurements of chatbot interactions, such as [10, 11, 12].
The architecture of HumAIne-chatbot integrates several key components, including an Interaction Tracking Engine, AI-Driven User Profiler, Prompt Manager, Dialogue History and Analysis Module, and Dialogue Recommendation Module. This comprehensive system enables the chatbot offer personalized interactions based on the individual user’s behavior and context. By leveraging personalized prompt enrichment, HumAIne-chatbot ensures that its responses are both relevant and contextually adapted to meet each user’s unique needs and capabilities.
Regarding the novel user profiling mechanism that created and utilizes virtual personas for LLMs by aggregating detailed backstories as described in [13]. This approach facilitates the training process of the AI user profiling approach, allowing the chatbot to generate personalized responses. By incorporating these virtual personas, HumAIne-chatbot can better align its conversational strategies with the broader context of the user’s values, beliefs, and objectives.
The remainder of the paper is structured as follows: Section 2 provides an overview of current conversational AI approaches, particularly focusing on personalization techniques and user profiling. Section 3 introduces the architecture of HumAIne-chatbot, detailing its components and the rationale behind integrating advanced user profiling mechanisms. Section 4 presents the plan for a user study, and Section 5 concludes the paper.
II Related Work
The development of human-centric chatbots is rooted in decades of research in AI, natural language processing (NLP), and human-computer interaction. This section provides an overview of the existing work that informs the design of HumAIne-chatbot, focusing on chatbot evolution, personalization techniques, and adaptive conversation management.
II-A Chatbot Evolution
The evolution of chatbots can be traced back to early systems such as ELIZA, developed in the 1960s, which employed basic pattern matching to simulate conversation [14].
With advancements in NLP, more sophisticated chatbots powered by LLMs, such as GPT-3 and GPT-4, have been introduced [5]. These LLMs enable the generation of coherent and contextually rich dialogues across various domains. However, despite these capabilities, LLM-based chatbots often provide generic responses that are not well adapted to the specific needs of individual users. ChatGPT, for example, is recognized as a leader in the field due to its public API, extensive training data, and adaptability across diverse tasks [15]. It has seen widespread adoption in sectors such as healthcare, education and finance [16], [17] [18] Moreover, MarketSense-AI is a notable application of GPT-4 in the financial domain, leveraging the Chain-of-Thought (CoT) approach to effectively elucidate investment decisions [19].
Another noteworthy application of LLMs, particularly the GPT-4 model, is presented by [20], which introduces a human-centric XAI tool named "x-[plAIn]". This tool represents a novel approach in Explainable Artificial Intelligence (XAI), aimed at making XAI more accessible to non-experts. It utilizes a custom LLM to generate clear, audience-specific summaries of XAI methods tailored to different groups. By adapting the explanations based on the audience’s knowledge level and interests, this model improves decision-making and accessibility.
HumAIne-chatbot is designed to be flexible, allowing it to integrate with any API-based LLM, providing adaptability and scalability in leveraging different models. This capability ensures that HumAIne-chatbot can incorporate the most suitable LLMs for specific use cases, thereby enhancing personalization and conversation quality.
II-B Personalization in Chatbots
Personalization is an important feature to improve user satisfaction in conversational systems. Existing frameworks, such as the user-adaptive model by [9], use explicit user input to tailor chatbot responses, requiring users to provide preferences and other personal data. This explicit personalization approach, while effective to some extent, presents privacy challenges and requires ongoing user effort. Most chatbot evaluation research has focused on technical metrics such as passing the Turing test, while there is a significant gap in studies that address user-centric aspects such as user satisfaction and motivation [8] [21]. Implicit personalization, which involves deriving user preferences from interaction behaviors, has been proposed to overcome these issues. [6] explored metrics like response time and typing speed to build user profiles, enabling adaptive conversation strategies. [22] emphasized the importance of holistic benchmarking frameworks that assess the quality, coherence, and adaptability of user interaction. However, these approaches still face limitations, particularly in effectively anticipating user needs or adjusting response depth in real-time for more nuanced and complex contexts.
Adaptive conversation management aims to make chatbot interactions more engaging and contextually relevant. Traditional scripted approaches often limit the chatbot’s ability to handle dynamic conversation paths. Deep reinforcement learning (RL) has been introduced to enable chatbots to learn optimal interaction policies dynamically [23].
HumAIne-chatbot goes beyond the state of the art by utilizing both explicit and implicit personalization methods. Additionally, we introduce an innovative approach for training the user-profiler based on [13], which creates detailed virtual personas from user backstories.
II-C Chatbots in Specific Domains
Chatbots are increasingly being utilized across a wide range of fields, from customer service to education, healthcare, and entertainment, serving both personal and professional purposes [24]. In customer service, chatbots like Dialogflow by Google and ChatGPT handle frequent customer inquiries and provide automated support. In healthcare, chatbots such as Woebot offer personalized support for mental health [25], demonstrating the value of tailored interaction.
In the restaurant industry, chatbots are being employed to enhance customer service processes and provide personalized recommendations [26]. These AI-powered assistants use advanced techniques such as embedding and context adaptation to generate accurate and relevant responses to customer queries. In the realm of elderly care, companion bots are gaining popularity as a means to combat loneliness and provide assistance with daily tasks [27]. These bots combine features of mobile robots and chatbots, offering functionalities such as medication reminders and entertainment options, thereby enhancing the quality of life for elderly individuals. The export business sector has also seen the integration of conversational agents to accelerate business processes [28]. These AI-powered chatbots use natural language processing and machine learning to facilitate real-time conversations, providing support in areas such as marketing and customer assistance. The banking sector has also adopted AI chatbots to improve customer loyalty [29]. Factors such as AI chatbot service quality, cognitive trust, and customer satisfaction play crucial roles in determining the effectiveness of these chatbots in banking applications.
III Proposed Methodology
In this section, we present our framework for personalizing chatbot interactions to improve user satisfaction. Our approach involves measuring various metrics that serve two main purposes: inputting data into the user profiler to personalize the HumAIne chatbot and evaluating the overall performance of the chatbot. The metrics used for personalization, as outlined in Table I, include both implicit and explicit measurements collected during interactions. Implicit metrics like session duration, response time, typing speed, grammatical accuracy, language complexity, and affect provide insights into the user’s behavior and communication style without requiring direct input. Explicit metrics, such as responses to pre-session questions, offer direct information from the user to inform the personalization process.
Metrics used to evaluate the overall performance of the chatbot, as shown in Table II, focus on assessing user satisfaction and the quality of the interaction. These include explicit metrics like user feedback ratings, comments, usability assessments, and measures of satisfaction such as expectation, impression, navigability, and engagement.
By combining these metrics, we aim to create a framework that not only personalizes the chatbot experience but also continually assesses and improves the chatbot’s performance.
TABLE I: Metrics for Personalization (Inputs to User Profiler)
Perspective Category Metrics Type Measurement
General Interaction Session Duration Length of Interaction Time Quantitative Implicit
Response Time Average Time to Respond Quantitative Implicit
Typing Speed Characters per Second Quantitative Implicit
Linguistic Grammatical Accuracy Frequency of Grammatical Errors Quantitative Implicit
Complexity of Language Avg. Sentence Length, Type-Token Ratio Quantitative Implicit
User Experience Affect Emotional Information, Personality Traits, Engagement Qualitative, Quantitative Implicit
General Interaction User Elicitation Responses to Pre-Session Questions Qualitative Explicit
TABLE II: Metrics for Evaluating Overall Performance of the Chatbot
Perspective Category Metrics Type Measurement
General Interaction User Feedback Ratings, Comments Qualitative Explicit
User Experience Usability Task Completion, Getting Assistance Qualitative, Quantitative Explicit
Satisfaction Expectation, Impression, Navigability, Engagement Qualitative, Quantitative Explicit
III-A Implicit Metrics
III-A1 Session Duration
Session Duration ($SD$) measures the length of time a user interacts with the chatbot in a single session. It can be calculated as:
$SD=t_{\text{end}}-t_{\text{start}}$ (1)
where $t_{\text{start}}$ is the timestamp when the session begins, and $t_{\text{end}}$ is when the session ends.
III-A2 Response Time
Response Time ($RT$) tracks how quickly the user responds to the chatbot’s prompts. For each turn $i$, the response time is:
$RT_{i}=t_{\text{user\_response}}^{i}-t_{\text{chatbot\_prompt}}^{i}$ (2)
The average response time over $N$ interactions is:
$\overline{RT}=\frac{1}{N}\sum_{i=1}^{N}RT_{i}$ (3)
III-A3 Sentiment Analysis
Sentiment Analysis involves quantifying the emotional tone of user responses. A sentiment score ($SS$) can be assigned to each user message using a sentiment analysis function $S(\cdot)$:
$SS_{i}=S(\text{User Message}_{i})$ (4)
The average sentiment score over $N$ messages is:
$\overline{SS}=\frac{1}{N}\sum_{i=1}^{N}SS_{i}$ (5)
[30]
III-A4 Grammatical Mistakes Frequency
Grammatical Mistakes Frequency ($GMF$) tracks the number of grammatical errors per word in the user’s responses. For message $i$:
$GMF_{i}=\frac{E_{i}}{W_{i}}$ (6)
where $E_{i}$ is the number of grammatical errors and $W_{i}$ is the total number of words in message $i$. The average over $N$ messages is:
$\overline{GMF}=\frac{1}{N}\sum_{i=1}^{N}GMF_{i}$ (7)
III-A5 Complexity of Language
The Complexity of Language ($CL$) can be quantified using metrics like average sentence length ($ASL$) and type-token ratio ($TTR$):
$CL_{i}=\alpha\times ASL_{i}+\beta\times TTR_{i}$ (8)
where $\alpha$ and $\beta$ are weighting factors, $ASL_{i}$ is the average number of words per sentence, and $TTR_{i}=\frac{\text{Number of Unique Words}}{\text{Total Words}}$ in message $i$.
III-A6 Typing/Interaction Speed
Typing Speed ($TS$) measures how quickly the user types their responses:
$TS_{i}=\frac{C_{i}}{t_{\text{user\_response}}^{i}-t_{\text{user\_start}}^{i}}$ (9)
where $C_{i}$ is the number of characters in message $i$, $t_{\text{user\_start}}^{i}$ is when the user starts typing, and $t_{\text{user\_response}}^{i}$ is when they send the message.
III-B Explicit Metrics
Explicit User Feedback can be quantified using a feedback score ($FS$) based on "likes" of the responses:
$FS=\frac{\sum_{i=1}^{n}L_{i}}{n}$ (10)
where $L_{i}$ represents the number of likes for response $i$, and $n$ is the total number of responses considered.
III-B1 Session-Based User Elicitation
To enhance user profiling before each session, we propose a novel approach where users are asked 2-3 targeted questions before starting a conversation. These questions are designed to gather updated user preferences, interests, or context-specific information that can enhance personalization during the interaction. This approach serves as an explicit metric for understanding the user’s needs at the beginning of each session and tailoring responses accordingly.
Asking these targeted questions before each session raises the issue of user elicitation, which involves finding the balance between asking relevant questions and not overwhelming the user. However, this elicitation process is not a straightforward task. Quantifying the performance of interview chatbots is a significant challenge. Developing a computational framework to measure the effectiveness of these chatbots is crucial for their improvement [31].
III-B2 Survey-Based Satisfaction
Survey-Based Satisfaction ($SBS$) aggregates responses from post-session surveys:
$SBS=\frac{1}{M}\sum_{j=1}^{M}R_{j}$ (11)
where $R_{j}$ is the user’s response to question $j$ on a standardized scale, and $M$ is the number of survey questions.
III-C Proposed Personalized Chatbot Architecture
Figure 1: Architecture of the HumAIne-chatbot system, illustrating the flow of data between components.
This section provides a detailed description of each component within the personalized chatbot system depicted in Figure 1. The components collectively enable the personalized and adaptive conversation flow, using advanced user profiling and recommendation strategies. While the sequence diagram 2 illustrates the interaction between the user and a RAG chatbot system, showcasing how the system retrieves relevant information, generates responses, and incorporates user feedback in an iterative process until the final response is approved.
III-C1 User Interface (UI) with Integrated Dialogue Management and Tracking
The User Interface UI) is the central point of interaction between the user and the chatbot system. It manages user input, displays responses, and handles interaction feedback, providing a seamless and engaging experience that ensures the conversation flow remains intuitive and accessible. The UI integrates both the Dialogue Manager and the Interaction Tracking Engine, allowing for cohesive management of the conversation and real-time adaptation based on user behavior.
The Dialogue Manager within the UI is responsible for controlling the overall flow of the conversation. It coordinates with other modules, such as the Dialogue History Module, AI-Driven User Profiler, and Prompt Manager, to maintain a coherent and adaptive conversational flow. By managing turn-taking, topic transitions, and conversation pacing, the Dialogue Manager ensures that the chatbot maintains a natural dialogue structure. It also integrates recommended prompts and manages user expectations, thereby enhancing the user experience.
The Interaction Tracking Engine embedded in the UI captures user activity during conversations. It continuously monitors all user interactions, including input characteristics such as response time, session duration and user behavior patterns. These metrics are then forwarded to other components for further analysis and personalization purposes.
III-C2 Personalization Metrics Collector
The Personalization Metrics Collector is responsible for gathering and analyzing metrics during user interactions, such as response times, engagement level, and other behavioral metrics. These metrics are critical for generating detailed user profiles, enabling adaptive personalization. The collected data feeds into the AI-Driven User Profiler, which refines user profiles for more precise personalization.
III-C3 Dialogue History and Analysis Module
The Dialogue History and Analysis Module is responsible for extracting and analyzing text-based metrics, including sentiment, grammatical accuracy, and complexity of language. This module also manages historical conversation data, including past user responses, previously discussed topics, and interaction patterns. The text-based metrics gathered by this module are forwarded to the AI-Driven User Profiler to enhance user profiling. Additionally, it plays a vital role in maintaining conversational continuity, allowing the chatbot to seamlessly reference past interactions and sustain context across the session.
Figure 2: Sequence Diagram of Interaction Flow of HumAIne-chatbot based on a RAG approach
III-C4 AI-Driven User Profiler with Reference Data Integration
The AI-Driven User Profiler is responsible for creating comprehensive user profiles using collected metrics, historical data, and reference data. It leverages ML models to analyze user interactions and extract information such as interests, and communication preferences. The integration of reference data provides a knowledge base against which user behavior and interactions are benchmarked. This reference data includes general user profiles created with the virtual personas for LLMs.
Initially, the AI model for user profiling will be trained as a supervised ML model using the gathered metrics and reference data. The supervised model will help identify relationships between user behavior metrics and corresponding user traits, forming the basis of user profiles. As the system matures, we will transition to a reinforcement learning (RL) model to further enhance personalization capabilities. The RL model will allow for continuous learning through interaction, optimizing the chatbot’s responses based on rewards derived from user engagement and satisfaction metrics.
To effectively utilize advanced techniques in personalized chatbots and create virtual personas for simulating chatbot interactions, we will incorporate the proposed metrics into our framework. This approach enables us to train an AI model for user profiling based on the gathered metrics from simulated behavior.
Phase I: Virtual Persona Pre-training
We leverage the Anthology of GPT-generated backstories to simulate a diverse set of virtual personas. Each backstory encodes rich demographic and behavioral traits (age, occupation, language style), conditioning an LLM to "play" a consistent user archetype [32]. By engaging these personas in multi-turn dialogues, we collect a large synthetic corpus with "ground-truth" profile labels. The profiler is first trained in a supervised manner on this corpus, learning to infer latent user attributes and preferences from conversational cues.
This approach aligns with the VirtualXAI framework, which uses GPT-generated personas for qualitative assessment of explainability methods [33].
Phase II: Online Reinforcement Learning Adaptation
In live deployment, each real user interaction constitutes an RL episode. The profiler agent receives a reward $r_{t}$ derived from engagement signals: conversational length, return sessions, and explicit feedback (likes, sentiment) [34, 35]. Formally, we define a reward function
$R(\tau)=\sum_{t=1}^{T}r_{t},\quad r_{t}=f(\text{turn\_length}_{t},\text{feedback}_{t}),$
where $\text{feedback}_{t}$ encapsulates sentiment and user ratings at turn $t$. This reward guides the policy $\pi(\mathbf{s}_{t}|\mathbf{h}_{t})$, mapping the user state $\mathbf{s}_{t}$ (interaction history) to profile updates and response strategies that maximize cumulative engagement.
Algorithmic Choice: Proximal Policy Optimization
We adopt Proximal Policy Optimization (PPO) [36] due to its stability and sample efficiency in high-dimensional, continuous action spaces. PPO updates policy parameters via clipped objective functions, ensuring gradual behavioral shifts—essential for maintaining user trust during personalization. The actor-critic architecture computes advantage estimates using a value network baseline, reducing variance in gradient estimates.
System Integration and Related Approaches
Our two-stage framework builds on recent work combining persona simulation and RL. Wang et al. [35] demonstrate efficacy of a supervised-to-RL pipeline for realistic user simulators, while Li et al. [34] show that policy gradient methods improve dialogue coherence and engagement.
III-D Prompt Manager
Based on the trained model, we will develop personalization strategies:
  1. 1.
    Dynamic Adaptation: Adjust the chatbot’s language complexity and response style based on the user’s inferred profile.
  2. 2.
    Topic Selection: Use the Dialogue Diversity (DD) metric to guide topic selection and maintain engaging conversations.
The Prompt Manager serves as the central component responsible for curating and managing conversation prompts. It receives input from the AI-Driven User Profiler to enrich the prompt based on the user’s profile and current context. The Prompt Manager consists of a extbfPrompt Curator submodule that refines and adjusts prompts to suit the user’s preferences and conversational style.
The user profile, constructed from the metrics in our JSON schemas, influences several key prompt parameters:
  • Language Complexity Level: Derived from the language_complexity metrics, this parameter adjusts the vocabulary richness, sentence structure complexity, and technical depth of the chatbot’s responses.
  • Response Detail Level: Based on engagement patterns and response times, this determines whether responses should be concise summaries or comprehensive explanations.
  • Domain-Specific Knowledge Calibration: Using grammar metrics and typing speed as proxies for expertise, this parameter tunes the level of domain-specific terminology and concepts included in responses.
  • Conversation Style: Sentiment analysis and feedback patterns inform the tone and formality of responses, ranging from professional to conversational.
The Dialogue Recommendation Module generates recommendations for the next conversational steps based on user inputs and current conversation dynamics. It uses insights from the User Profiler and Interaction Tracking Engine to decide which follow-up questions or suggestions would best suit the user’s current state. This module is essential for maintaining user engagement, offering proactive and personalized suggestions that align with user interests, and facilitating smooth transitions between topics.
III-D1 LLM/API/RAG System and Overall Flow
The LLM/API/RAG System forms the core response generation mechanism for HumAIne-Chatbot, which includes large language models (LLMs), APIs, and retrieval-augmented generation (RAG) mechanisms. The enriched prompts generated by the Prompt Manager are forwarded to this system to generate highly contextual and informative responses. Depending on the user’s query, it can access external knowledge via APIs or utilize vector databases to retrieve relevant information. Importantly, our approach is independent of the underlying LLM technology that supports HumAIne-Chatbot, making it adaptable to different LLMs or retrieval mechanisms as required. By designing the system to be independent of specific LLM technologies, HumAIne-Chatbot remains flexible and capable of leveraging ongoing advancements in language models, retrieval methods, and other supporting technologies. This flexibility ensures that as newer, more advanced LLMs or retrieval systems are developed, they can be seamlessly integrated to enhance the chatbot’s capabilities without requiring fundamental changes to the system architecture.
IV Evaluation Strategy
To validate HumAIne-chatbot, we adopted a virtual persona–based evaluation approach inspired by VirtualXAI framework [37] that has used the approached proposed by [13]. Instead of human recruitment, we generated 50 diverse virtual personas with systematically varied demographics (18–65+, balanced across education and occupations), communication preferences (language complexity, detail level, style), expertise levels (beginner–expert across finance, health, education, technology), behavioral patterns (patience, engagement, multitasking), and emotional profiles (stress, confidence, response speed). This design ensures scalability, reproducibility, diversity without recruitment bias, and ethical efficiency (privacy, consent, fatigue).
We performed a controlled A/B experiment contrasting two systems: (i) a Personalized HumAIne-chatbot using an AI-driven profiler with online adaptation; and (ii) a Baseline Non-Personalized Chatbot identical in all components but without adaptive profiling. Each persona interacted with both systems over 10 topic domains (e.g., career development, personal finance, health, technology, environmental sustainability). The study produced 150 sessions (3 per persona), averaging 23 messages and 4.13 minutes per session (SD = 0.16), with a 100% completion rate and response efficiency of 334.55 messages/min.
The metrics framework focused primarily on user satisfaction (1–5 Likert, positive feedback ratio, engagement ratings, helpfulness). Complementary metrics assessed session efficiency (duration, turns to completion, task completion) and personalization effectiveness (response relevance, detail-level match, language complexity adaptation, style consistency). Statistical analysis employed descriptive stats, t-tests (and Mann–Whitney U), ANOVA for multi-factor comparisons, and effect sizes (Cohen’s $d$). Post-hoc power analysis confirmed adequate power to detect medium–large effects.
V Results
The persona cohort exhibited balanced demographics: ages 18–25 (14%), 26–35 (16%), 36–45 (20%), 46–55 (16%), 56–65 (24%), 65+ (10%); education levels: High School (20%), Some College (18%), Bachelor’s (8%), Master’s (26%), PhD (18%), Professional Certification (10%). The diversity index was 0.13. Conversations showed 100% completion, 23.0 messages/session on average, 4.13 minutes/session (SD = 0.16), and throughput of 334.55 messages/min.
0.05–0.100.10–0.150.15–0.200.20–0.250.25–0.300.30–0.35022446688101012121414Satisfaction Score RangeNumber of SessionsDistribution of Satisfaction Scores Across All Sessions
Figure 3: Distribution of satisfaction scores across all 50 conversation sessions.
18-2526-3536-4546-5556-6565+022446688101012121414Age GroupNumber of PersonasAge Distribution
HSSCBAMAPhDPC022446688101012121414Education LevelNumber of PersonasEducation Distribution
Figure 4: Demographic distribution of virtual personas: (a) Age groups, (b) Education levels.
Satisfaction was the central metric. The overall mean satisfaction was 0.173 (median = 0.182, SD = 0.071). The experimental (personalized) condition outperformed control by +45.0% (0.173 vs. 0.119), significant at $p<0.001$ (t = 4.394; Mann–Whitney U = 1791.0, $p<0.001$).
TABLE III: Descriptive Statistics for Control and Experimental Groups
Group Mean Std Dev Median Range
Control (Non-Personalized) 0.119 0.050 0.121 0.034–0.246
Experimental (Personalized) 0.173 0.071 0.183 0.056–0.318
Difference +0.054 +0.021 +0.062 +0.072
TABLE IV: Primary Outcome Results
Outcome Control Experimental Improvement
Mean Satisfaction 0.119 0.173 +45.0%
95% CI [0.105, 0.133] [0.153, 0.193]
TABLE V: Secondary Outcome Measures
Metric Control Experimental Difference
Relevance Score 0.152 0.198 +0.046
Personalization Score 0.089 0.234 +0.145
Expertise Alignment 0.167 0.189 +0.022
Style Match 0.134 0.187 +0.053
Task Achievement 0.142 0.198 +0.056
Across topic domains, improvements were consistent, with the largest gains in professional networking (+50.3%) and creative projects (+50.9%).
TABLE VI: Subgroup Analysis by Topic Domain
Topic Control Experimental Improvement
Professional Networking 0.156 0.235 +50.3%
Creative Projects 0.127 0.192 +50.9%
Career Development 0.083 0.124 +48.8%
Education and Learning 0.100 0.149 +48.5%
Environmental Sustainability 0.135 0.199 +48.1%
Personal Finance 0.138 0.200 +44.4%
Technology Trends 0.094 0.134 +42.8%
Health and Wellness 0.095 0.133 +39.5%
Work-Life Balance 0.126 0.176 +39.5%
Travel and Culture 0.134 0.184 +36.8%
Overall 0.119 0.173 +45.0%
We also observed cross-session learning: profile accuracy improved over 2–3 sessions, with clearer gains in detail-level matching and language adaptation; variability persisted in complex multi-domain scenarios.
The evaluation methodology offered several important strengths. It provided systematic control and ensured reproducibility, while supporting multi-metric coverage and statistical rigor. The framework was also highly scalable and carried notable ethical advantages, as no human data collection was required. At the same time, certain limitations must be acknowledged. The synthetic nature of virtual personas cannot fully capture the complexity of real human behavior, and the structured scenarios used in the study may not reflect the nuances of natural conversation. Moreover, the evaluation was limited to short-term sessions, leaving open questions about how the system would generalize to long-horizon, free-form interactions.
VI Conclusions
The evaluation of HumAIne-chatbot demonstrated the value of the VirtualXAI approach as a scalable, reproducible, and ethically responsible methodology for assessing personalization in conversational systems. By leveraging virtual personas, the framework enables systematic exploration of user diversity while avoiding the practical and ethical constraints of human-subject studies. Looking ahead, future work will focus on enriching profiling with multimodal signals, refining reinforcement learning objectives to capture long-term user goals, and conducting pilot deployments to validate the system’s effectiveness in real-world settings.
VII Acknowledgements
The research leading to the results presented in this paper has received funding from the Europeans Union’s funded Project HumAIne under grant agreement no 101120218.